# FACT-GS: Frequency-Aligned Complexity-Aware Texture Reparameterization for 2D Gaussian Splatting

### v2

@@ 5.4 Runtime and Efficiency
As shown in Table 3, FACT-GS maintains real-time rendering and remains within $1\text{–}2$ FPS of Textured GS [chao2025textured] on NeRF Synthetic [mildenhall2021nerf] using a single RTX A6000 GPU. The sampling-density alignment is implemented as a per-primitive CUDA warp computed in parallel, keeping the rasterization pipeline unchanged and thus introducing negligible overhead.
@@ 5.5 Ablation and Analysis
Figure 5: Per-Gaussian average RGB gradient magnitude. Textured GS [huang2024textured] concentrates most textures in the low-frequency range (<10<10), indicating insufficient capacity for high-frequency regions. Our method reallocates capacity according to local signal complexity, yielding more high-frequency textures under the same parameter budget.
Texture Frequency Analysis. To quantify the effect of our reparameterization, we measure the spatial frequency of each learned per-Gaussian RGB texture $\mathbf{T}\in\mathbb{R}^{\tau\times\tau\times 3}$ using the mean gradient magnitude [eskicioglu2002image]:
@@
Fig. 5 shows the resulting distributions. Visually, Textured GS [huang2024textured] forms a narrow peak in the low-frequency range (<10<10), indicating that its uniform grid under-samples visually complex regions. In contrast, FACT-GS produces a broader and right-shifted distribution, reflecting substantially more high-frequency textures under the same parameter budget.
@@
Ablation on Texture Parameters. Fig. 6 evaluates reconstruction quality as the number of per-Gaussian texture parameters increases from $100$ to $3600$, using $10\%$ of the default 2DGS primitives. Across all parameter budgets, FACT-GS achieves higher PSNR than Textured GS [huang2024textured]. Notably, FACT-GS matches the reconstruction quality of Textured GS while using only $\tfrac{1}{16}$ of its texture parameters. These results confirm that frequency-aligned sampling enhances the efficiency of texture space utilization compared to uniform-grid textures, offering substantially greater improvements than simply increasing the texture resolution.

### v3

@@ 5.4 Runtime and Efficiency
As shown in Table 3, FACT-GS maintains real-time rendering and remains within $1\text{–}2$ FPS of Textured GS [4] on NeRF Synthetic [25] using a single RTX A6000 GPU. The sampling-density alignment is implemented as a per-primitive CUDA warp computed in parallel, keeping the rasterization pipeline unchanged and thus introducing negligible overhead.
@@ 5.5 Ablation and Analysis
Figure 5: Per-Gaussian average RGB gradient magnitude. Textured GS [15] concentrates most textures in the low-frequency range (<10<10), indicating insufficient capacity for high-frequency regions. Our method reallocates capacity according to local signal complexity, yielding more high-frequency textures under the same parameter budget.
Texture Frequency Analysis. To quantify the effect of our reparameterization, we measure the spatial frequency of each learned per-Gaussian RGB texture $\mathbf{T}\in\mathbb{R}^{\tau\times\tau\times 3}$ using the mean gradient magnitude [5]:
@@
Fig. 5 shows the resulting distributions. Visually, Textured GS [15] forms a narrow peak in the low-frequency range (<10<10), indicating that its uniform grid under-samples visually complex regions. In contrast, FACT-GS produces a broader and right-shifted distribution, reflecting substantially more high-frequency textures under the same parameter budget.
@@
Ablation on Texture Parameters. Fig. 6 evaluates reconstruction quality as the number of per-Gaussian texture parameters increases from $100$ to $3600$, using $10\%$ of the default 2DGS primitives. Across all parameter budgets, FACT-GS achieves higher PSNR than Textured GS [15]. Notably, FACT-GS matches the reconstruction quality of Textured GS while using only $\tfrac{1}{16}$ of its texture parameters. These results confirm that frequency-aligned sampling enhances the efficiency of texture space utilization compared to uniform-grid textures, offering substantially greater improvements than simply increasing the texture resolution.
@@ Acknowledgment
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), under funding reference numbers RGPIN-2021-03477.
Supplementary Material
7 Appendix
7.1 Author Contributions
@@
7.2 Datasets
We evaluate our method on five standard benchmarks for novel view synthesis, covering both synthetic and real-world scenes. The experimental scenes are listed below:
  • NeRF Synthetic [25]: all scenes.
  • MipNeRF 360 v2 [3]: publicly available scenes including Bicycle, Counter, Garden, Room, Bonsai, Kitchen, and Stump.
  • DTU [16]: Scan105, Scan110, Scan37, and Scan63.
  • Tanks & Temples [18]: Train, Truck, and Church.
  • LLFF [26]: Horns and Fortress.
7.3 Budgets for Fig. 3 and Fig. LABEL:fig:teaser
  • NeRF Ship: $10\%$.
  • NeRF Chair: $10\%$
  • NeRF Hotdog: $1\%$
  • Tanks & Temples Truck: $1\%$.
  • Tanks & Temples Train: $1\%$
  • Tanks & Temples Church: $10\%$
  • MipNeRF 360 Kitchen: $10\%$.
7.4 Ablation on Texture Parameters
[…]
Across all parameter budgets and datasets, FACT-GS achieves higher PSNR and SSIM, and lower LPIPS than Textured GS [15], which demonstrates that our method improves the efficiency of texture space utilization compared with uniform-grid textures, yielding substantially greater performance gains than simply increasing the texture resolution across datasets.
Parameters Count. To ensure a fair comparison with Textured GS, we matched the parameter number of both methods as closely as possible; nevertheless, our method consistently uses fewer parameters. The texture parameters of Textured GS are computed as $\tau_{\text{tex}}\times\tau_{\text{tex}}\times 4$, while those of our method are $\tau_{\text{FACT}}\times\tau_{\text{FACT}}\times 6$. We report the exact number of texture parameters for both methods in Table 4.
7.5 Model Size
Model size is matched for TexturedGS and our method in the main paper with identical Gaussian counts and total texture parameter budgets; an explicit size-matched comparison with 2DGS is further reported in Tab. 5 and Fig. 9.
7.6 Interpreting RGB Gradient.
We provide additional qualitative evidence for texture capacity utilization in Fig. 7. RGB gradient magnitude provides a practical proxy for local spatial frequency and effective texel allocation under a fixed texel budget, as higher gradients correspond to more rapidly varying color signals. Fig. 7(C–D) visualizes per-Gaussian local texture footprints, where our method concentrates color variation into a more compact and structured support, producing sharper transitions and clearer edges than TexturedGS. Accordingly, Fig.5 shows a right-shift in the RGB gradient distribution under parameter-matched settings, indicating increased effective utilization of texture capacity for high-frequency detail reconstruction.
Figure 7: RGB Gradient and per-Gaussian Texture Visualization.
[…]
Textured GS Ours
$5\times 5\times 4=100$ $4\times 4\times 6=96$
$10\times 10\times 4=400$ $8\times 8\times 6=384$
$15\times 15\times 4=900$ $12\times 12\times 6=864$
$20\times 20\times 4=1600$ $16\times 16\times 6=1536$
$30\times 30\times 4=3600$ $24\times 24\times 6=3456$
Table 4: Exact texture parameters count of Textured GS and our method for the ablation results in Fig. 8.
Number of Primitives. We report the average number of primitives in the default 100% 2DGS setting in Table 6.
Method Average Size Tanks &Temples LLFF DTU
(MB) PSNR ($\uparrow$)  /  SSIM ($\uparrow$)  /  LPIPS ($\downarrow$)
2DGS 100.94 22.79 / 0.805 / 0.194 27.68 / 0.881 / 0.113 27.85 / 0.905 / 0.133
Ours 99.23 23.11 / 0.816 / 0.177 28.46 / 0.891 / 0.106 28.90 / 0.910 / 0.128
Table 5: Size-matched comparison with 2DGS.
Figure 9: Quantitative size-matched comparison with 2DGS.
7.7 Backpropagation of FACT Texture
Forward. Let $\mathbf{T}_{i}\in\mathbb{R}^{\tau\times\tau\times 4}$ denote the texture map of the $i$-th Gaussian, and $\mathbf{D}_{i}\in\mathbb{R}^{\tau\times\tau\times 2}$ its learned deformation field. Given a camera ray from pixel $\mathbf{p}$ that intersects the $i$-th Gaussian at local coordinates $(u,v)$, the sampled displacement $(\Delta u,\Delta v)$ is computed by bilinear interpolation over $\mathbf{D}_{i}$:
$\Delta u=\sum_{j\in N_{(u,v)}}b_{j}(u,v)\,D_{j}^{u},\hskip 28.80008pt\Delta v=\sum_{j\in N_{(u,v)}}b_{j}(u,v)\,D_{j}^{v},$ (10)
where $N_{(u,v)}$ denotes the four neighboring texels surrounding $(u,v)$, and $b_{j}(u,v)$ are their corresponding bilinear interpolation weights. Let $\Phi(u,v)=(u,v)+(\Delta u,\Delta v)$ be the frequency-aligned sampling coordinate. The sampled color $c$ from texture $\mathbf{T}_{i}$ is then obtained as
$c=\sum_{k\in N_{\Phi(u,v)}}b_{k}(\Phi(u,v))\,T_{k},$ (11)
where $N_{\Phi(u,v)}$ are the four neighboring texels around $\Phi(u,v)$ and $b_{k}(\Phi(u,v))$ are the corresponding interpolation weights.
Backward. The gradient of the sampled color $c$ with respect to $D_{j}^{u}$ at texel $j$ is given by
$\frac{\partial c}{\partial D_{j}^{u}}=\Big(\sum_{k\in N_{\Phi(u,v)}}T_{k}\,\frac{\partial b_{k}(\Phi(u,v))}{\partial(u+\Delta u)}\Big)\frac{\partial\Delta u}{\partial D_{j}^{u}}.$ (12)
Substituting Eq. (10) into Eq. (12) yields
$\frac{\partial c}{\partial D_{j}^{u}}=\Big(\sum_{k\in N_{\Phi(u,v)}}T_{k}\,\frac{\partial b_{k}(\Phi(u,v))}{\partial(u+\Delta u)}\Big)b_{j}(u,v),$ (13)
which defines the differentiable gradient flow from the rendered color $c$ to the deformation field $\mathbf{D}_{i}$.
Number of primitives
NeRF Synthetics [25] 91,396
MipNeRF 360 v2 [3] 2,763,229
DTU [16] 316,989
Tanks & Temple [18] 1,584,660
LLFF [26] 490,511
Table 6: The average number of primitives under the default 100% optimization setting.
Methods NeRF Synthetic [25] MipNeRF 360v2 [3] DTU [16] Tanks & Temples [18] LLFF [26]
[…]
Textured GS (100) 30.88 0.949 0.0542 24.38 0.675 0.3710 28.72 0.899 0.1640 22.65 0.792 0.2206 28.29 0.884 0.1206
Ours (96) 31.51 0.954 0.0439 24.84 0.703 0.3247 28.80 0.904 0.1504 22.83 0.802 0.2036 28.46 0.891 0.1058
Textured GS (400) 31.33 0.953 0.0457 24.72 0.695 0.3242 28.46 0.898 0.1543 22.80 0.801 0.1974 28.36 0.885 0.1084
Ours (384) 31.89 0.957 0.0364 25.25 0.728 0.2710 28.51 0.901 0.1431 23.00 0.810 0.1791 28.68 0.896 0.0952
Textured GS (900) 31.49 0.954 0.0411 24.82 0.702 0.3000 28.31 0.898 0.1513 22.84 0.803 0.1847 28.17 0.886 0.1075
Ours (864) 31.90 0.957 0.0341 25.31 0.734 0.2506 28.61 0.901 0.1382 23.05 0.813 0.1684 28.60 0.894 0.0932
Textured GS (1600) 31.55 0.954 0.0383 24.85 0.707 0.2850 28.39 0.897 0.1475 22.86 0.805 0.1760 28.09 0.886 0.1037
Ours (1536) 31.87 0.957 0.0331 25.31 0.733 0.2420 28.48 0.900 0.1409 23.06 0.814 0.1619 28.49 0.894 0.0900
Textured GS (3600) 31.52 0.954 0.0359 24.82 0.707 0.2705 27.26 0.897 0.1441 22.84 0.804 0.1669 27.85 0.883 0.1013
Ours (3456) 31.73 0.956 0.0326 25.25 0.731 0.2374 28.44 0.898 0.1383 22.96 0.811 0.1567 28.32 0.889 0.0960
[…]